OpenStudy (anonymous):

Help solve using quadratic equation: y=5x^2+5x-5

OpenStudy (anonymous):

please help...

OpenStudy (anonymous):

the quadratic formula is: \[x = (-b \pm \sqrt{b^2 - 4ac}) / 2a\] looking at your polynomial, you can tell that a = 5, b = 5, and c = -5 so plug that into the formula.

jhonyy9 (jhonyy9):

so there are 5x^2 +5x -5 =y ,than for you solve it make equal zero,like 5x^2 +5x -5 =0 divide both sides by 5 and will get x^2 +x -1 =0 discriminant D=1+4=5 so than those roots x_1 and x_2 =(-1+/- sqrt5)/2 =(1+/-sqrt5)/2 x_1 = (1-sqrt5)/2 x_2 =(1+sqrt5)/2 hope so much that is understandably sure ,right good luck bye